# A Multi-Task and Multi-Label Classification Model for Implicit Discourse Relation Recognition

### v2

@@ Abstract
We address the inherent ambiguity in Implicit Discourse Relation Recognition (IDRR) by introducing a novel multi-task classification model capable of learning both multi-label and single-label representations of discourse relations. Our model is trained exclusively on the DiscoGeM corpus and evaluated both on the DiscoGeM and the PDTB 3.0 corpus. We establish the first benchmark on multi-label IDRR classification and achieve SOTA results on single-label IDRR classification using the DiscoGeM corpus. Finally, we present the first evaluation on the potential of transfer learning between the DiscoGeM and the PDTB 3.0 corpus on single-label IDRR classification.
@@ 1 Introduction
Discourse analysis is responsible for understanding the relations of coherence between segments of a text. While different frameworks have been proposed for the analysis of discourse, one of the most widely used is the Penn Discourse Treebank (PDTB) (Miltsakaki et al., 2004; Prasad et al., 2008). In the PDTB framework, a discourse relation is established between two textual segments (called arguments) either explicitly, through the use of a discourse connective, or implicitly when there is no connective. The sense of the relation is then determined from a list of 36 senses organized hierarchically in three levels (Prasad et al., 2019). The first level has four broad senses and the following levels have subsequently more detailed senses. At the implicit level, the recognition of the sense of the relation - known as implicit discourse relation recognition (IDRR) - remains arguably the hardest task in discourse analysis. This comes as a result of the inherently subjective nature of the task, where even trained human annotators find it difficult to agree on the sense annotation of implicit discourse relations (Rohde et al., 2016; Scholman and Demberg, 2017; Hoek et al., 2021).
In fact, such ambiguity is present in most tasks where language is concerned, since different language constructs can carry different meanings or interpretations (Poesio, 2020; Basile et al., 2021; Jiang and de Marneffe, 2022). Yet, most NLP classification tasks are established on the idea that a single gold-label should exist and neglect that there could be genuine human variation in labeling due to disagreement, subjective interpretation or the existence of multiple plausible answers (Pavlick and Kwiatkowski, 2019; Uma et al., 2021; Plank, 2022). In the field of discourse, the idea that discourse annotation can often be ambiguous is not new (Stede, 2008) and has been recently highlighted by the difficulties in establishing a clear mapping between implicit discourse relations across frameworks (Demberg et al., 2019; Costa et al., 2023).
For this reason, a growing body of work has recently considered disagreement between humans annotators not as noise, but as an opportunity to enhance the accuracy and relevance of NLP models by capturing the complexity of human interpretations and biases (Nie et al., 2020; Uma et al., 2020; Fornaciari et al., 2021; van der Meer et al., 2024). This new perspective has lead to a recent paradigm shift in implicit discourse annotation. For instance, rather than relying on expert annotators to find a single gold label for each implicit relation, the DiscoGeM corpus (Scholman et al., 2022a) crowdsourced this task to multiple laypeople to obtain a distribution of labels for each relation following the PDTB 3.0 framework (Prasad et al., 2019).
This paper addresses the inherent ambiguity in IDRR by proposing novel multi-task implicit discourse relation classification model capable of learning both multi-label and single-label representations of discourse relations. Our model is trained exclusively on the DiscoGeM corpus (Scholman et al., 2022a) and evaluated both on the DiscoGeM and the PDTB 3.0 corpus (Prasad et al., 2019). Overall, the main contributions of this work are:
@@
  • We present the first fully multi-label approach to IDRR by training a multi-task classification model on the DiscoGeM corpus to produce probability distributions across multiple labels for all instances. Our model can then be further adapted to the traditional single-label task by selecting the label with the highest probability from the generated distribution.
@@
  • We establish the first benchmark on multi-label IDRR classification and achieve SOTA results on single-label IDRR classification using the DiscoGeM corpus.
@@
  • We present an in-depth analysis on the potential of transfer learning between the DiscoGeM and the PDTB 3.0 corpus on single-label IDRR classification.
@@ 2 Previous Work
The advent of transformer-based models (Vaswani et al., 2017) has lead to a significant improvement in performance across all NLP tasks and in IDRR in particular. While traditional model would rely on recurrent neural networks (Shi and Demberg, 2019a; Liang et al., 2020), recent models leverage the power of pre-trained large language models (LLMs) through fine-tuning (Long and Webber, 2022; Liu and Strube, 2023; Wu et al., 2023) or prompt-tuning (Zhou et al., 2022; Zhao et al., 2023; Chan et al., 2023) to achieve SOTA results. However, the SOTA performance of these models has only slightly increased over the last years.
One possible explanation for this lack of significant improvement can be attributed to the fact that current approaches to IDRR rely on the principle that there is a single ground truth in the interpretation of implicit discourse relations (Aroyo and Welty, 2013, 2015; Dumitrache et al., 2019). However, IDRR is an inherently ambiguous task, often with possible multiple interpretations (Rohde et al., 2016; Scholman and Demberg, 2017; Hoek et al., 2021). Thus, a new line of work has recently tried to capture this source of ambiguity by pioneering new discourse annotated corpora capturing a distribution of sense labels per relation instead of a single reference label (Yung et al., 2019; Pyatkin et al., 2020; Scholman et al., 2022a, b; Pyatkin et al., 2023). Following this line of work, recent works have leveraged the potential of such multi-labelled discourse relations for IDRR (Yung et al., 2022; Pyatkin et al., 2023; Costa and Kosseim, 2024). However, even if trained on multi-labelled data, these models still predict a single-label sense per relation. On the other hand, Long et al. (2024) leveraged the $4.9\%$ of implicit discourse relations in the PDTB 3.0 (Prasad et al., 2019) annotated with two senses to develop classification models capable of predicting up to two senses per relation. Yet, to the best of our knowledge, no previous work has truly captured the full multi-label distribution of all possible senses across all instances.
@@ 3 Our Approach
In this work we propose a novel multi-task classification model capable of simultaneously predicting probability distributions across all three sense levels in the PDTB 3.0 framework. Figure 1 shows the architecture of our model. The model takes as input a concatenated pair of discourse arguments (ARG1+ARG2) and generates an embedding using the pre-trained RoBERTa (Liu et al., 2019) language model [^1] We also experimented with other pre-trained models, but RoBERTa achieved the best performance (see Section 5.1).. The embedding is then fed to a linear and dropout layer, before being passed to three separate classification heads - one for each of the three sense levels. In the multi-label setting, the output of each classification head is a probability distribution over the existing senses at the corresponding level. In the single-label setting, an additional pooling layer selects the sense with the highest probability in the corresponding distribution as the output of each classification head. We use the Adam optimization algorithm (Kingma and Ba, 2015) to minimize the loss function of our model, which we calculate as the sum of the individual losses of each classification head.
Figure 1: The model takes as input a pair of discourse arguments and generates one output for each sense level in the PDTB 3.0 framework. In the multi-label setting each output is a vector of probabilities over the existing senses at each level, while in the single-label setting each output is the sense (selected by the pooling layers in faded colors) with the highest probability score.
@@
To calculate the loss of each multi-label classification head, we use the mean absolute error (MAE) loss function (see Equation 2 in Appendix B) as it leads better performance (see Section 5.1). Similarly to other multi-label classification works in NLP (Pyatkin et al., 2023; van der Meer et al., 2024), we evaluate our model in the multi-label classification setting using the Jensen-Shannon (JS) distance (Lin, 1991) to measure the similarity between the predicted and the target probability distributions.
@@
To calculate the loss of each single-label head, we used the cross-entropy (CE) loss function (see Equation 1 in Appendix B) as it emphasizes the correct classification of the highest probability sense label (see Section 5.1). To evaluate our model in the single-label classification setting, we use the weighted F1-score between the labels with the highest scores in the predicted and the target distributions.
@@ 4 Data Preparation
We used two different discourse annotated corpora: DiscoGeM (Scholman et al., 2022a) and the PDTB 3.0 (Prasad et al., 2019). Our model was trained exclusively on the DiscoGeM corpus and evaluated on both DiscoGeM (for multi- and single-label classification) and on the PDTB 3.0 (for single-label classification). Although they use different annotation methodologies, both corpora follow the same discourse framework (Webber et al., 2019). This enabled our model to leverage the knowledge acquired from DiscoGeM and directly apply it to the PDTB 3.0 through transfer learning.
@@ 4.1 DiscoGeM
The DiscoGeM corpus (Scholman et al., 2022a) contains 6,807 inter-sentential implicit discourse relations distributed across four textual genres: 2,800 relations from political texts, 3,060 from literary texts, 645 from encyclopedic texts and 302 relations extracted from the PDTB 3.0 corpus (Prasad et al., 2019). Each relation was annotated by at least 10 different laypeople. These annotations were then averaged to produce a multi-label sense distribution across 29 different senses in the PDTB 3.0 framework - the Belief and the Speech-Act level-2 senses were not annotated in DiscoGeM.
To use DiscoGeM, we adapted the standard set of 14 level-2 senses proposed by Kim et al. (2020) for IDRR classification in the PDTB 3.0. Since DiscoGeM did not include the Cause+Belief sense from the standard set, and to maintain a 14 label set, we replaced it by the Similarity sense as it was the most frequent of the remaining level-2 senses [^2] We could not replace Cause+Belief by another Contingency sense, because they were all already considered.. In addition to this set of level-2 senses, we also considered level-3 senses when possible. When no level-3 sense was available, we selected the corresponding level-2 sense as replacement. Table 8 in Appendix A shows the distribution of senses across all levels in the original DiscoGeM and in our adapted set of senses.
Figure 2: Distribution of majority label senses across all splits of the DiscoGeM corpus.
For single-label classification, we replace the full multi-label sense distribution of each discourse relation in DiscoGeM by its sense with the highest score - referred to as the majority label. To reduce computational complexity and maintain consistency across model evaluations, we used the same training, validation and testing splits in both the multi and single-label settings. We split 70% of the corpus for training, 10% for validation and 20% for testing. To ensure a balanced distribution, we kept the same distribution of majority labels across all data splits as shown in Figure 2. We opted for a fixed training, validation and test split, instead of multiple fold splits for cross-validation, to streamline the experimentation process as we conducted an extensive number of experiments.
@@
The majority of research in IDRR relies on the single-label annotated PDTB 3.0 corpus (Prasad et al., 2019). The PDTB 3.0 corpus was annotated by expert annotators and consists of 53,631 discourse relations extracted from Wall Street Journal news articles - from which 21,827 are implicit. Since there are currently no other benchmarks in multi-label IDRR, we also evaluated our model on the traditional single-label classification task using different test splits of the PDTB 3.0 corpus.
To allow transfer learning between DiscoGeM and the PDTB 3.0, we used the same set of 14 level-2 senses described in Section 4.1. Following the common approach to IDRR, with the PDTB 3.0 we only kept the level-1 and level-2 senses. To compare our work against SOTA models in single-label IDRR, we generated the two commonly used Lin (Lin et al., 2009) and Ji (Ji and Eisenstein, 2015) test splits - the former uses section 23 in the PDTB 3.0 corpus as the test set, while the latter uses sections 21-22. However, since recent works highlighted the limitations of using such small test sets to draw meaningful conclusions (Shi and Demberg, 2017; Kim et al., 2020), we also generated test splits following the cross-validation scheme proposed by Kim et al. (2020). Table 9 in Appendix A shows the total number of level-2 sense instances on the different single-label test sets.
@@ 4.3 Corpora Agreement
Since 302 discourse relations from the PDTB 3.0 were co-annotated in DiscoGeM, we calculated the annotation agreement between both corpora. Table 1 shows the number of relations where the reference label from the PDTB 3.0 agreed with the majority label (top-1) of DiscoGeM, or was found among the top-3, top-5 or top-10 labels in DiscoGeM. The fact that for 87 (18.8%) of the jointly annotated relations, the reference label from the PDTB 3.0 corpus was not selected by at least one of the annotators of DiscoGeM highlights the inherent ambiguity in IDRR.
@@ 5 Experiments
In order to optimize performance in terms of lower JS distances for the multi-label classification and in terms of higher weighted F1-scores for the single-label classification, we evaluated different configurations of our model (see Fig. 1) on the DiscoGeM validation set (see Fig. 2). All of the reported results are the average scores of three different runs with different random starts for 10 epochs and with a batch size of 16. We ran all of our experiments on a 32-core compute node with 512GB of RAM and all of the code used can be found on GitHub [^3] https://anonymous-link.
5.1 Model Selection
LM Loss JS Distance (Multi-Label) F1-Score (Single-Label)
@@ 5.1 Model Selection
Starting with the claim that LMs pre-trained with a next sentence prediction (NSP) objective lead to better results in IDRR (Shi and Demberg, 2019b), we evaluate the performance of using BERT (Devlin et al., 2019) as the backbone of our model against RoBERTa (Liu et al., 2019), which was used in recent works on single-label IDRR (Long and Webber, 2022; Liu and Strube, 2023) and was not pre-trained on NSP. We then experimented with distilled versions of both models (Sanh et al., 2019). Finally, we evaluated the performance of each LM with respect to the MAE and the CE loss functions as shown in Table 2.
Table 2 shows that for both loss functions the RoBERTa model outperforms BERT both in terms of lower JS distances in the multi-label setting and higher F1-scores in the single-label setting across all sense levels. This shows that the benefit of using larger models (RoBERTa against BERT) outweighs previous findings that LMs pre-trained with the NSP training objective lead to better results in IDRR (Shi and Demberg, 2019b). Following this observation, both RoBERTa and BERT models outperform their distilled versions across all metrics. Finally, Table 2 also shows that the choice of loss function plays an important role in multi- and single-label IDRR. Training the model with the MAE loss leads to lower JS distances in all configurations, while training the model with the CE loss leads to better F1-scores. Therefore, we chose RoBERTa as the backbone of the model and use the MAE loss when training for multi-label IDRR and the CE loss when training for single-label IDRR. We also experimented with the mean squared error (MSE) and the Huber loss. However, they performed slightly worse as shown in Table 10 in Appendix B.

### v3

@@ Abstract
We propose a novel multi-label classification approach to implicit discourse relation recognition (IDRR). Our approach features a multi-task model that jointly learns multi-label representations of implicit discourse relations across all three sense levels in the PDTB 3.0 framework. The model can also be adapted to the traditional single-label IDRR setting by selecting the sense with the highest probability in the multi-label representation. We conduct extensive experiments to identify optimal model configurations and loss functions in both settings. Our approach establishes the first benchmark for multi-label IDRR and achieves SOTA results on single-label IDRR using DiscoGeM. Finally, we evaluate our model on the PDTB 3.0 corpus in the single-label setting, presenting the first analysis of transfer learning between the DiscoGeM and PDTB 3.0 corpora for IDRR.
@@ 1 Introduction
Implicit discourse relation recognition (IDRR) is one of the most challenging tasks in computational discourse analysis. Its goal is to identify the sense of discourse relations connecting text arguments in the absence of explicit discourse connectives, such as but and because. The Penn Discourse Treebank (PDTB) (Miltsakaki et al., 2004; Prasad et al., 2008a) organizes discourse senses across three hierarchical levels with increasing degrees of detail (Webber et al., 2019). However, understanding the sense of a discourse relation can be a complex and subjective task without the guidance of discourse connectives. Consider the two arguments of the implicit discourse relation below:
Arg1: The lights flickered.
Arg2: The power went out.
In this example, it is not clear whether the two arguments of the relation are related by a temporal or causality relation. One could interpret that the lights flicked and then the power went out, or that the lights flickered because the power went out. This ambiguity illustrates the biggest challenge of IDRR. Despite its difficulty, IDRR plays a crucial role in downstream tasks that rely on text coherence. In summarization, for example, it helps preserve the logical flow and rhetorical intent of the source text. In dialogue systems, it supports coherence and intent modeling by inferring how utterances relate to each other, improving dialogue discourse parsing (Li et al., 2024; Thompson et al., 2024).
To date, most research on IDRR has relied on the different iterations of the PDTB corpus (Prasad et al., 2008b, 2019). However, despite significant efforts, SOTA performance on IDRR has plateaued in recent years at F1-Scores of $71.59$ at level-1 and $57.62$ at level-2 (Long and Webber, 2022; Liu and Strube, 2023; Chan et al., 2023; Zhao et al., 2023; Zeng et al., 2024; Long and Webber, 2024). One possible reason may be the inherently subjective nature of implicit discourse interpretation, which can be difficult to capture using predominantly single-label annotated corpora (Stede, 2008; Rohde et al., 2016; Scholman and Demberg, 2017; Hoek et al., 2021). The idea of possible multiple relations between discourse arguments had already been considered in the Segmented Discourse Representation Theory (SDRT) (Asher and Lascarides, 2003) and the ambiguity in the annotation of implicit relations has been further highlighted by the challenges in mapping them across discourse frameworks (Demberg et al., 2019; Costa et al., 2023). In light of these challenges, recent studies have advocated for the multi-label annotation of implicit relations to better capture their nuanced and complex nature in discourse corpora (Yung et al., 2019; Pyatkin et al., 2020; Scholman et al., 2022a, b; Pyatkin et al., 2023; Yung and Demberg, 2025). This shift in perspective has led to the creation of DiscoGeM (Scholman et al., 2022a) - the first multi-label annotated corpus of implicit discourse relations.
Following recent work in NLP (Pavlick and Kwiatkowski, 2019; Nie et al., 2020; Basile et al., 2021; Fornaciari et al., 2021; Uma et al., 2021; Plank, 2022; Jiang and de Marneffe, 2022), we see disagreement between humans annotators not as noise, but as an opportunity to enhance the relevance of IDRR models by capturing the complexity of human interpretations. This paper addresses the subjectivity in IDRR by proposing a multi-task classification model capable of learning both multi-label and single-label representations of implicit discourse relations. Our model is trained on the DiscoGeM corpus and uses a multi-task architecture to leverage the interdependence between senses and jointly learn sense representations across all three sense levels in the PDTB 3.0 hierarchy. The main contributions of this work are:
@@
  • We present the first multi-label approach in IDRR which produces probability distributions over all possible sense labels of a discourse relation using the DiscoGeM corpus. Our multi-label model can then be adapted to the traditional single-label task by selecting the label with the highest probability.
  • We conduct extensive experiments, comparing different pre-trained language models as encoders of our model and evaluating how different loss functions impact performance in multi-label and single-label IDRR.
@@
  • We establish the first benchmark on multi-label IDRR and achieve SOTA results on single-label IDRR using DiscoGeM.
@@
  • We present an in-depth analysis on the potential of transfer learning between the DiscoGeM and the PDTB 3.0 corpora on single-label IDRR.
@@ 2 Previous Work
[…]
The shift toward multi-label annotation has stirred an initial wave of research in multi-label IDRR. For instance, Long et al. (2024) used the 4.9% of implicit relations in the PDTB 3.0 corpus that are annotated with two senses to build a model capable of predicting up to two senses per instance. However, given that the 95.1% of PDTB 3.0 annotations remain single-label, their model predominantly produces single-label predictions. In contrast, Yung et al. (2022) and Costa and Kosseim (2024) also use the multi-label DiscoGeM corpus but convert its annotations into a single-label format during training. Thus, to date, no previous work has truly captured the full multi-label representation of implicit discourse relations - with the exception of our work on multi-lingual IDRR (Costa and Kosseim, 2025), where we consider multi-label IDRR in a multi-lingual setting incorporating hierarchical learning in the training of our classification model and comparing it against LLMs via direct prompting with few-shot learning. Similar to other works (Chan et al., 2024; Yung et al., 2024), we found that prompting LLMs leads to worst results compared to the fine-tuning of PLMs such as RoBERTa in the context of IDRR. Therefore, we do not consider prompting LLMs in this paper.
@@ 3 Our Approach
In this work, we introduce a novel multi-task classification model for IDRR that simultaneously predicts sense distributions across all three levels in the PDTB 3.0 framework. As illustrated in Figure 1, the model processes a concatenated pair of discourse arguments (ARG1+ARG2) using RoBERTa (Liu et al., 2019) as the PLM encoder [^1] We also experimented with other PLMs, but RoBERTa achieved the best performance (see Section 5.1).. The resulting contextualized embedding is passed through a linear transformation and dropout layer and subsequently fed to three distinct classification heads - each corresponding to a sense level in the hierarchy. In the multi-label setting, each head outputs a probability distribution over the available senses at its respective level. In the single-label setting, we apply an additional pooling layer that selects the sense with the highest probability from each distribution. We use the Adam optimizer (Kingma and Ba, 2015) to minimize the loss function, which we calculate as the sum of the individual losses of each classification head.
ARG1ARG2+RoBERTa-baseLinear + Dropout LayerClass Head Level-1Linear LayerNorm. LayerPool. LayerL1 Sense(s)Class Head Level-2Linear LayerNorm. LayerPool. LayerL2 Sense(s)Class Head Level-3Linear LayerNorm. LayerPool. LayerL3 Sense(s)
Figure 1: Architecture of our multi-task classification model for IDRR. Given a pair of discourse arguments as input, the model generates one output per sense level in the PDTB 3.0 framework. In the multi-label setting, each output is a probability distribution over the senses at the corresponding level. In the single-label setting, the sense with the highest probability is selected via the pooling layers (indicated by dashed lines).
@@
For each multi-label classification head, we compute the loss using the mean absolute error (MAE) loss function (see Equation 2 in Appendix B), which we found to yield better performance (as detailed in Section 5.1). Following previous work in multi-label classification for NLP (Pyatkin et al., 2023; van der Meer et al., 2024), we evaluate model performance in this setting using the Jensen-Shannon (JS) distance (Lin, 1991) to measure the similarity between the predicted and target probability distributions.
@@
In the single-label setting, we compute the loss for each classification head using the cross-entropy (CE) loss function (see Equation 1 in Appendix B), which emphasizes the correct classification of the highest probability sense label (as discussed in Section 5.1). We evaluate model performance in this setting using the weighted F1-score based on the majority label in the predicted and target sense distributions.
@@ 4 Data Preparation
We used two different discourse annotated corpora in this work: DiscoGeM (Scholman et al., 2022a) and PDTB 3.0 (Prasad et al., 2019). Our model was trained exclusively on the DiscoGeM corpus and evaluated on both DiscoGeM (for multi-label and single-label classification) and on PDTB 3.0 (for single-label classification). Although the two corpora differ in annotation methodologies, both follow the same discourse framework (Webber et al., 2019). This shared foundation allowed us to directly evaluate the transfer learning potential of training our model on DiscoGeM and evaluating it on PDTB 3.0 in single-label IDDR.
@@ 4.1 DiscoGeM
The DiscoGeM corpus contains 6,807 implicit discourse relations drawn from four textual genres: 2,800 relations from political texts, 3,060 from literary texts, 645 from encyclopedic texts and 302 relations extracted from the PDTB 3.0 corpus. Each relation was independently annotated by at least 10 crowdworkers and the resulting annotations were aggregated to form a probability distribution over 29 discourse senses in the PDTB 3.0 framework - the Belief and the Speech-Act level-2 senses were not annotated in DiscoGeM.
Since the DiscoGeM corpus did not include the Cause+Belief sense from the standard set of 14 level-2 senses proposed by Kim et al. (2020) for IDRR classification in the PDTB 3.0 framework, we replaced it with the Similarity sense (the most frequent among the remaining level-2 senses) to preserve a 14-label set. In addition to classifying senses at level-1 and level-2, we also incorporated level-3 sense labels when available. When no level-3 sense was available, we defaulted to their corresponding level-2 sense as replacement. Table 8 in Appendix A shows the distribution of senses across all levels in the original DiscoGeM annotations and in our adapted label set.
For single-label classification, we replaced the multi-label sense distribution of each discourse relation by the sense with the highest score - referred to as the majority label. To reduce computational complexity and maintain consistency across model evaluations, we used the same training, validation and testing splits in both the multi-label and the single-label settings. We split 70% of the corpus for training, 10% for validation and 20% for testing. To ensure a balanced distribution, we kept the same distribution of majority labels across all data splits (see Figure 2 in Appendix A). We opted for a fixed split of the data, instead of multiple folds for cross-validation, to streamline the experimentation process as we conducted an extensive number of experiments.
@@
The PDTB 3.0 corpus was annotated by expert annotators and consists of 53,631 discourse relations extracted from Wall Street Journal news articles - from which 21,827 are implicit. Since there are currently no other benchmarks in multi-label IDRR using DiscoGeM and the majority of research in IDRR relies on the single-label annotated PDTB 3.0 corpus, we also evaluated our model on the traditional single-label classification task using different test splits of the PDTB 3.0 corpus through transfer learning.
To allow transfer learning between the two corpora, we used the same set of 14 level-2 senses described in Section 4.1. Following the common approach to IDRR, we only kept the level-1 and level-2 senses in the PDTB 3.0. To compare our work against SOTA models in single-label IDRR, we replicated the two commonly used Lin (Lin et al., 2009) and Ji (Ji and Eisenstein, 2015) test splits - the former uses section 23 in the PDTB 3.0 corpus as the test set, while the latter uses sections 21-22. However, since recent works highlighted the limitations of using such small test sets to draw meaningful conclusions (Shi and Demberg, 2017; Kim et al., 2020), we also generated test splits following the cross-validation scheme proposed by Kim et al. (2020). Table 9 in Appendix A shows the total number of level-2 sense instances on the different single-label test sets.
@@ 4.3 Corpora Agreement
Since a total of 302 implicit discourse relations taken from the PDTB 3.0 were also independently annotated in DiscoGeM, we calculated the annotation agreement between the two corpora over this overlapping set of relations. Table 1 reports the number of relations where the PDTB 3.0 reference label matches the majority label (top-1) in DiscoGeM, or appears within its top-3, top-5, or top-10 majority labels. The fact that for 28.8% of the jointly annotated relations, the reference label from the PDTB 3.0 corpus was not selected by at least one of the annotators of DiscoGeM illustrates the variability in human interpretation of implicit discourse relations - even when annotations are based on the same underlying framework.
@@ 5 Experiments
To optimize our model (illustrated in Figure 1), we explored multiple model configurations and loss functions using the validation set of DiscoGeM. All of the results reported in the following sections are the average scores of three different runs with different random starts for 10 epochs and with a batch size of 16. We ran our experiments on a 32-core compute node with 512GB of RAM. All of the code used can be found on GitHub [^2] https://github.com/nelsonfilipecosta/Implicit-Discourse-Relation-Recognition.
PLM Loss JS Distance
$\searrow$
(Multi-Label)
F1-Score
$\nearrow$
(Single-Label)
@@
Table 2: Results of experimenting with different pre-trained language models (PLMs) and different loss functions in multi-label classification (JS distance) and in single-label classification (weighted F1-score). The results were averaged across three different runs with random starts. Values in bold show the best score for each metric.
5.1 Model Selection
Motivated by prior work suggesting that PLMs trained with a next sentence prediction (NSP) objective perform better on IDRR (Shi and Demberg, 2019), we compared BERT (Devlin et al., 2019), which includes the NSP objective, against RoBERTa (Liu et al., 2019), which omits it. RoBERTa has also been widely adopted in SOTA single-label IDRR models (Long and Webber, 2022; Liu and Strube, 2023; Zhao et al., 2023; Zeng et al., 2024; Long and Webber, 2024), making it a strong baseline. To further explore efficiency-performance trade-offs, we also evaluated distilled variants of both models (Sanh et al., 2019). Each PLM configuration was evaluated using both the mean absolute error (MAE) and cross-entropy (CE) loss functions. The results of these experiments are summarized in Table 2.
As shown in Table 2, RoBERTa consistently outperforms BERT across both loss functions, yielding lower Jensen-Shannon (JS) distances in the multi-label setting and higher F1-scores in the single-label setting across all sense levels. These results suggest that the performance gains of using larger models outweigh the previously reported advantages of using models pre-trained on the NSP objective in IDRR (Shi and Demberg, 2019). In addition, both BERT and RoBERTa outperform their respective distilled variants, reaffirming the benefit of using full-sized models for this task. The results in Table 2 also show that the choice of loss function plays an important role. Models trained with MAE consistently achieve lower JS distances in the multi-label setting, while model trained with CE achieve higher F1-scores in the single-label setting. We present a more detailed analysis on the performance of each loss function in Appendix B, as well as the results of training the different models with the mean squared error (MSE) and the Huber loss - which led to slightly worse results as shown in Table 10 of the appendix. Based on these findings, we selected RoBERTa as the backbone encoder of our model, using MAE loss for multi-label classification and CE loss for single-label classification.